### v1

# On the Fundamental Limit of Stochastic Gradient Identification Algorithm Under Non-Persistent Excitation

Abstract
Stochastic gradient methods are of fundamental importance in system identification and machine learning, enabling online parameter estimation for large-scale and data-streaming processes. The stochastic gradient algorithm stands as a classical identification method that has been extensively studied for decades. Under non-persistent excitation, the best known convergence result requires the condition number of the Fisher information matrix to satisfy $\kappa(\sum_{i=1}^{n}\varphi_{i}\varphi_{i}^{\top})=O((\log r_{n})^{\alpha})$, where $r_{n}=1+\sum_{i=1}^{n}\|\varphi_{i}\|^{2}$, with strong consistency guaranteed for $\alpha\leq 1/3$ but known to fail for $\alpha>1$. This paper establishes that strong consistency in fact holds for the entire range $0\leq\alpha<1$, achieved through a novel algebraic framework that yields substantially sharper matrix norm bounds. Our result nearly resolves the four-decade-old conjecture of chen1986limit, bridging the theoretical gap from $\alpha\leq 1/3$ to nearly the entire feasible range.
keywords:
System identification, stochastic gradient algorithm, strong consistency, non-persistent excitation, fundamental limit
thanks: This research was supported by the National Natural Science Foundation of China under grant No. 12288201 and the National Key R&D Program of China under grant 2024YFC3307201.
These authors contributed equally to this work.
1 Introduction
How can we develop mathematical models for real-world physical processes from noisy observational data to support critical engineering tasks such as controller design, prediction, and fault diagnosis? This question lies at the heart of system identification, see LJUNG20101. In the era of artificial intelligence, characterized by large-scale data and online learning requirements, stochastic gradient (SG) algorithms have regained prominence. These methods support iterative, online processing of randomly perturbed data while ensuring convergent parameter estimation. They have also become the foundation for numerous algorithmic variants, including Adam, see ADAM.
The SG method is closely related to stochastic approximation (SA), a framework originating in the pioneering work of robbins1951stochastic. Subsequent milestones include the stochastic optimization approach of kiefer1952stochastic, the general convergence theorem of dvoretzky1956stochastic, asymptotic analyses by chung1954stochastic and sacks1958asymptotic, and the ODE method of gladyshev1965stochastic, which together established a rigorous mathematical foundation for this class of stochastic iterative algorithms.
The application of this theory to system identification emerged in the 1960s. astrom1965numerical established a systematic offline identification framework using maximum likelihood estimation, introducing the pivotal concept of persistent excitation (PE). Concurrently, sakrison1964application pioneered the application of stochastic approximation to online identification, shifting the focus from offline modeling to real-time estimation. During the 1970s, research evolved toward closed-loop systems. The groundbreaking work of astrom1972self initiated the theoretical analysis of stochastic adaptive control, while ljung1977analysis advanced the ODE method of gladyshev1965stochastic to provide a unified tool for analyzing the asymptotic behavior of general stochastic algorithms.
Although SG algorithms possess well-established strong consistency under PE conditions, see ljung1977analysis; anderson1979strong; chen1981strong, their convergence behavior in the absence of such conditions warrants critical investigation. One motivation comes from adaptive control, see goodwin1980discrete, where a decoupling between parameter estimation and optimal control is observed: rigorously enforcing PE ensures strong consistency but may not yield an optimal controller, see chen1984strong; conversely, pursuing an optimal controller can compromise parameter consistency, see becker1985adaptive. A second practical challenge stems from the fact that many systems operate under insufficiently rich inputs, as empirically observed in diverse domains such as quadrotor UAVs, see chowdhary2012experimental, neural networks, see nar2019persistencyexcitationrobustnessneural, and spherical parallel robots, see RAD2020104026. Without theoretical guarantees in such scenarios, inadequate excitation can lead to severely biased parameter estimates, unacceptably slow convergence, or even outright divergence of the identification algorithm.
The seminal works of chen1982strong and lai1982 independently and concurrently broke new ground by systematically weakening PE condition required for the strong consistency of LS. chen1982strong established that strong consistency holds if the condition number of the Fisher information matrix grows as
$\kappa\left(\sum_{i=1}^{n}\phi_{i}\phi_{i}^{\top}\right)=O(r_{n}^{\delta})\quad\mathrm{a.s.},\quad 0\leq\delta<\frac{1}{2}$ (1)
where $r_{n}=1+\sum_{i=1}^{n}\|\phi_{i}\|^{2}$. Simultaneously, Lai and Wei, working within a general stochastic regression framework, derived that strong consistency holds if
$\kappa\left(\sum_{i=1}^{n}\phi_{i}\phi_{i}^{\top}\right)=o\left(\frac{r_{n}}{\log r_{n}}\right)\quad\mathrm{a.s.},$ (2)
Building upon this perspective, the Chen-Guo approach offers a profound understanding of SG algorithm under relaxed excitation. chen1985strong determines the convergence speed of SG algorithm in this non-PE case. Through a remarkable sample-path-wise analysis that relates strong consistency to the convergence of an instrumental matrix sequence $\Phi(n,0)$, see chen1985consistency, under mild conditions on the noise (not necessarily i.i.d. or of bounded variance), chen1985strong; guo1993; chen1986limit showed that if
$\kappa\left(\sum_{i=1}^{n}\phi_{i}\phi_{i}^{\top}\right)=O((\log r_{n})^{\alpha})\quad\mathrm{a.s.},$ (3)
then:
  • $\alpha\leq 1/3$ ensures strong consistency.
  • $\alpha>1$ precludes strong consistency.
This stark dichotomy inevitably raises a fundamental question: What are the fundamental limits of SG identification algorithms? To what extent can we rely on the results of SG algorithms? More broadly, it challenges us to understand how much insight into a system can be gained from finite, noisy, and poorly-structured observations in a world replete with uncertainties.
chen1986limit conjectured that the established sufficient condition is inherently conservative. They posited that strong consistency should, in fact, be guaranteed for the entire range $\alpha\leq 1$, with $\alpha=1$ representing the ultimate “critical excitation” boundary for the algorithm. For decades, the validity of this conjecture—and thus the characterization of the algorithm’s fundamental limits within the gap $1/3\leq\alpha\leq 1$ —has remained an open problem, representing a significant challenge in the theoretical foundations of SG identification.
The main contribution of this paper is a nearly complete resolution of this long-standing conjecture. Our results bridge the theoretical gap as follows:
  • We prove that for the entire range $0\leq\alpha<1$, the condition (3) is indeed sufficient for the almost sure convergence of the stochastic gradient algorithm to the true parameters, thereby providing a near-complete characterization of its behavioral landscape.
  • To establish this result, we develop a novel algebraic approach that offers a more transparent and versatile framework for obtaining the requisite matrix norm bounds, superseding previous analytical obstacles.
The remainder of this paper is organized as follows: Section 2 revisits the problem formulation and Chen-Guo approach. Section 3 details our new algebraic approach bounding matrix products. Section 4 presents detailed integral estimates. Section 5 proves the main theorem, and Section 6 offers concluding remarks.
Notation: Throughout this paper, the logarithm $\log$ denotes the natural logarithm with base $e$. $A_{n}\lesssim B_{n}$ implies that there exists a positive constant $C$ such that $A_{n}\leq CB_{n}$. The Landau notation $O(\cdot)$ and $\Theta(\cdot)$ follow their standard asymptotic definitions. $\|\cdot\|$ denotes the Euclidean norm for vectors and the spectral norm for matrices, unless specified otherwise. $\|\cdot\|_{F}$ means the Frobenius norm. For a nonsingular matrix $A$, the condition number $\kappa(A)$ is defined by $\kappa(A)=\|A\|\|A^{-1}\|$.
2 Problem Formulation
In this section, we reformulate the problem and briefy review Chen-Guo approach. Consider the MIMO system
$y_{n}+A_{1}y_{n-1}$ $+\cdots+A_{p}y_{n-p}$ (4)
$=B_{1}u_{n-1}+\cdots+B_{q}u_{n-q}+\varepsilon_{n},$
where $y_{n}$ and $u_{n}$ denote the $d$-dimensional output and the $l$-dimensional input respectively, $A_{i},i=1,\ldots,p,B_{j},j=1,\ldots,q$ are the unknown matrices to be identified, $\varepsilon_{n}$ is the $d$-dimensional system noise driven by a martingale difference sequence $\{w_{n}\}$, that is,
$\varepsilon_{n}=w_{n}+C_{1}w_{n-1}+\cdots+C_{r}w_{n-r},$ (5)
and $w_{n}$ satisfies
$\mathbb{E}(w_{n}|\mathcal{F}_{n-1})=0,\quad\forall n\geq 1,$ (6)
where $\{\mathcal{F}_{n}\}$ is a family of nondecreasing $\sigma$-algebras defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ and $C_{k},k=1,2,\ldots,r$ are unknown matrices.
Let $z$ be the shift-back operator, define
$C(z)=I+C_{1}z+\cdots C_{r}z^{r},$ (7)
and set
$\theta^{\top}=\begin{bmatrix}-A_{1}&\cdots&-A_{p}&B_{1}&\cdots&B_{q}&C_{1}&\cdots&C_{r}\end{bmatrix}$, (8)
$\varphi_{n}^{\top}=[y_{n-1}^{\top},\cdots,y_{n-p+1}^{\top},$ (9)
$u_{n-1}^{\top},\cdots,$ $u_{n-q+1}^{\top},$
$\hat{w}_{n-1}^{\top},\cdots,\hat{w}_{n-r+1}^{\top}],$
where we use the priori estimate of $w_{n}$, which is also called the estimated residuals:
$\hat{w}_{n-1}:=y_{n}-\theta_{n-1}^{\top}\varphi_{n-1}.$ (10)
It is clear that $\varphi_{n}$ is defined recursively, which is $\mathcal{F}_{n}$-measurable. The estimation problem is cast as a linear regression problem:
$y_{n+1}=\theta^{\top}\varphi_{n}+w_{n+1}.$ (11)
Denote by $\theta_{n}$ the estimate for $\theta$ at time $n$. Given any determined initial values $\theta_{0}$ and $\varphi_{0}$, we can state the SG algorithm as follows:
$\theta_{n+1}=\theta_{n}+\frac{\varphi_{n}}{r_{n}}(y_{n+1}^{\top}-\varphi_{n}^{\top}\theta_{n}),$ (12)
where $r_{n}$ is given by
$r_{n}=1+\sum_{i=1}^{n}\|\varphi_{i}\|^{2},\quad r_{0}=1.$ (13)
Chen and Guo introduce the following instrumental transition matrix:
$\Phi(n+1,i)=(I-A_{n})\Phi(n,i),\quad n\geq i,$ (14)
$\quad\Phi(i,i)=I,$ (15)
where we denote
$A_{n}=\frac{\varphi_{n}\varphi_{n}^{\top}}{r_{n}}.$ (16)
Chen and Guo leverage the strictly positive real (SPR) condition to bound the error introduced by the estimated residuals, thereby ensuring that the parameter error dynamics vanish as the instrumental transition matrix converges to zero.
Theorem 1 (Chen and Guo, 1985)
If $r=0$, or $r>0$ but $C(z)-\frac{1}{2}I$ is SPR, then $\Phi(n,0)\to 0$ implies $\theta_{n}\to\theta$.
Notably, under mild noise assumptions, this sufficient condition is also necessary when $r=0$.
Assumption 1 (“Condition A”)
  • As $n\to\infty$, $\sum_{i=0}^{n}\frac{\varphi_{i}}{r_{i}}\varepsilon_{i+1}^{\top}$ tends to a finite limit $S$.
  • There exist $c>0$ and $\delta>0$ which may depend on $\omega$ such that $\left\|S-\sum_{i=0}^{n-1}\frac{\varphi_{i}}{r_{i}}\varepsilon_{i+1}^{\top}\right\|\leq cr_{n}^{-\delta}$.
Remark 1
“Condition A” imposes a constraint on the cumulative noise effect, requiring that the weighted sum of noise terms converges sufficiently fast. This condition is weaker than typical independence or stationarity assumptions and allows for correlated noise sequences, making it applicable to a wide range of practical scenarios. The parameter $\delta$ quantifies the rate of this convergence.
Theorem 2 (Chen and Guo, 1985)
Assume $r=0$ and the noise sequence $\{\varepsilon_{n}\}$ satisfy “Condition A” at a sample $\omega\in\Omega$, then for this given sample path, for any initial value $\theta_{0}$, we have $\theta_{n}\to\theta$ if and only if $\Phi(n,0)\to 0$, and in this case, we have convergence speed
$\|\theta_{n}-\theta\|=O(\|\Phi(n,0)\|^{\delta/1+\delta}),$ (17)
where $\delta>0$ possibly depends on the sample $\omega$.
This transforms the problem of establishing strong consistency for the SG algorithm into one of proving that $\Phi(n,0)\to 0$. In essence, this is a problem of bounding the norm of a deterministic matrix product.
Assumption 2
$w_{n}$ is a $\mathcal{F}_{n}$-measurable process satisfying
  • $\mathbb{E}[w_{n}\mid\mathcal{F}_{n-1}]=0$.
  • $\mathbb{E}[\|w_{n}\|^{2}\mid\mathcal{F}_{n-1}]\leq c_{0}r_{n-1}^{\varepsilon}$, where $c_{0}>0$, $0\leq\varepsilon\leq 1$.
Remark 2
The bound $r_{n-1}^{\varepsilon}$ allows for potential growth in the noise variance, with $\varepsilon=0$ corresponding to the bounded variance case and $\varepsilon>0$ permitting variance that grows with the accumulated regressor norm $r_{n}$. Assumption 2 can often be used to verify that “Condition A” holds.
Assumption 3
$r_{n}\to\infty$, and $r_{n}=O(r_{n-1})$.
Remark 3
The condition $r_{n}\to\infty$ ensures we have enough information for identification. In real-world applications, most physical systems exhibit bounded input-output behavior due to physical constraints, actuator limits, and sensor ranges. The condition $r_{n}=O(r_{n-1})$ naturally arises in such scenarios, as it implies that the energy injected into the system cannot grow arbitrarily fast between consecutive time steps.
Using analytical techniques, Chen and Guo were able to significantly relax the excitation requirements for strong consistency of parameter estimates and provided quantitative convergence rates.
Theorem 3 (Guo, 1993)
Provided that “Condition A”, Assumption 3 and condition number growth
$\kappa\left(\sum_{i=1}^{n}\varphi_{i}\varphi_{i}^{\top}\right)=O((\log r_{n})^{1/3})$ (18)
are satisfied at a sample $\omega\in\Omega^{\prime}$, then for this given sample path, we have $\Phi(n,0)\to 0$. Moreover, we have convergence speed
$\|\theta_{n}-\theta\|=O((\log r_{n})^{-\delta}),$ (19)
where $\delta>0$ possibly depends on the sample $\omega$.
Remark 4
This theorem establishes that the logarithmic growth of the condition number is sufficient for strong consistency of the SG algorithm. The condition number bound $O((\log r_{n})^{1/3})$ represents a remarkably weak excitation condition, significantly relaxing earlier requirements of persistent excitation. The convergence rate $O((\log r_{n})^{-\delta})$ shows that parameter estimates converge almost surely, though the rate can be slow depending on the sample path.
Through a counterexample inspired by lai1982, Chen and Guo show that if the condition number grows strictly faster than $O(\log r_{n})$, the SG algorithm can be made to fail.
Theorem 4 (Chen and Guo, 1986)
Assume Assumption 2 is satisfied almost surely, then for any $\delta>0$, there exists a random vector sequence $\{\varphi_{n}\}$ satisfies Assumption 3 and the condition number growth
$\kappa\left(\sum_{i=1}^{n}\varphi_{i}\varphi_{i}^{\top}\right)=O((\log r_{n})^{1+\delta}),\ \mathrm{a.s.}$ (20)
but $\Phi(n,0)\not\to 0,\mathrm{a.s.}$
Chen and Guo posit that the established threshold for the condition number’s growth rate might not be the final word, suggesting that the critical exponent in the bound could be sharpened to its ultimate limit $O(\log r_{n})$.
3 Mathematical Framework
In this section, to ensure our theoretical framework is sufficiently general to address a broad class of problems, we let $0\leq A_{n}\leq I$ denote general positive semi-definite symmetric matrices, which admit a decomposition of the form
$A_{n}=\phi_{n}\phi_{n}^{\top}.$ (21)
Our goal is to estimate products involving the instrumental matrix (Theorem 5). We introduce an auxiliary sequence:
$x_{i+1}=(I-A_{i})x_{i},\quad i\geq k.$ (22)
From this, we obtain
$x_{i}-x_{k}=\sum_{j=k}^{i-1}A_{j}x_{j},$ (23)
and since $A_{i}^{2}\leq A_{i}$, it follows that
$\|x_{i}\|^{2}\leq\|x_{i-1}\|^{2}-\langle A_{i}x_{i-1},x_{i-1}\rangle.$ (24)
Summing over $j$ yields
$\sum_{j=k}^{i-1}\|\phi_{j}^{\top}x_{j}\|^{2}\leq\|x_{k}\|^{2}-\|x_{i}\|^{2}.$ (25)
Now, introduce a nonnegative real sequence $\mu_{n}\geq 0$ called the weights and define a weighted sum $S_{ik}$ of the matrices $A_{n}$ over the interval $[k,i]$:
$S_{ik}=\sum_{j=k}^{i-1}\mu_{j}A_{j}.$ (26)
Remark 5
The design of $S_{ik}$ represents one of the fundamental innovations in this framework. The weights $\mu_{j}$ serve multiple purposes: they can compensate for non-uniform regressor magnitudes, emphasize periods of high information content, or discount older measurements in time-varying systems.
We estimate the quadratic form:
$x_{k}^{\top}S_{ik}x_{k}$ $=x_{k}^{\top}\left(\sum_{j=k}^{i-1}\mu_{j}A_{j}\right)x_{k}=\sum_{j=k}^{i-1}\mu_{j}x_{k}^{\top}A_{j}x_{k}$ (27)
$=\sum_{j=k}^{i-1}\mu_{j}x_{k}^{\top}\phi_{j}\phi_{j}^{\top}x_{k}=\sum_{j=k}^{i-1}\mu_{j}\|\phi_{j}^{\top}x_{k}\|^{2}.$
To gain better insight into this quantity, define the vectors
$v=\begin{bmatrix}\ldots&\phi_{j}^{\top}x_{k}&\ldots\end{bmatrix}^{\top}$, (28)
$u=\begin{bmatrix}\ldots&\phi_{j}^{\top}x_{j}&\ldots\end{bmatrix}^{\top}$. (29)
Note that
$u-v=\begin{bmatrix}\ldots&\phi_{j}^{\top}(x_{j}-x_{k})&\ldots\end{bmatrix}^{\top}$. (30)
Multiplying both sides of (23) by $\phi_{i}^{\top}$ gives
$(u-v)_{j}=\phi_{j}^{\top}x_{j}-\phi_{j}^{\top}x_{k}=\sum_{l=k}^{j-1}(\phi_{j}^{\top}\phi_{l})(\phi_{l}^{\top}x_{l}).$ (31)
Since $\phi_{l}^{\top}x_{l}=u_{l}$, we define a strictly lower-triangular matrix $C$ by
$C_{jl}=\phi_{j}^{\top}\phi_{l},\quad k\leq l<j.$ (32)
By construction, we have the matrix identity:
v=(IC)uv=(I-C)u (33)
Remark 6
The matrix $C$ encodes the inter-temporal correlation structure of the regressor sequence, i.e. it quantifies how much information each new regressor $\phi_{i}$ shares with previous regressors $\phi_{l},l<j$. The strictly lower-triangular structure reflects the causal nature of time: future regressors cannot influence past ones. The matrix identity reveals that the initial projection error $v_{j}$ equals the concurrent projection error $u_{j}$ minus a correction term that accounts for how much the state has evolved due to previous updates.
Set $\Lambda=\mathrm{diag}(\ldots,\sqrt{\mu_{j}},\ldots)$, then
$\lambda_{\min}(S_{ik})\|x_{k}\|^{2}$ $\leq x_{k}^{\top}S_{ik}x_{k}=\|\Lambda v\|^{2}$ (34)
$=\|\Lambda(I-C)u\|^{2}\leq\|\Lambda(I-C)\|^{2}\|u\|^{2}$
$=\|\Lambda(I-C)\|^{2}(\|x_{k}\|^{2}-\|x_{i}\|^{2}).$
It is a standard fact that the operator norm of a matrix can be bounded by its Frobenius norm:
$\frac{1}{\sqrt{\mathrm{rank}(A)}}\|A\|_{F}\leq\|A\|\leq\|A\|_{F}.$ (35)
This yields a good estimate of the matrix norm:
$\|\Lambda(I-C)\|$ $\leq\|\Lambda\|+\|\Lambda C\|\leq\|\Lambda\|+\|\Lambda C\|_{F}$ (36)
$=\sqrt{\max_{k\leq j<i}\mu_{j}}+\sqrt{\sum_{j=k}^{i-1}\mu_{j}\sum_{l=k}^{j-1}(\phi_{j}^{\top}\phi_{l})^{2}}.$
Remark 7
The separate estimation of $\Lambda$ and $\Lambda C$ is strategically important, because it allows us to distinguish between two fundamentally different sources of “complexity” in the system: the amplitude of weights (controlled by $\|\Lambda\|$) and the temporal correlation structure (captured by $\|\Lambda C\|$). In applications, this means we can independently control system behavior through weight selection (choice of $\mu_{j}$) and through regressor design (affecting correlation structure $C$).
For convenience, we denote
$B_{jk}=\sum_{l=k}^{j-1}(\phi_{j}^{\top}\phi_{l})^{2}.$ (37)
Thus, we obtain the following inequality relating $\|x_{k}\|$ and $\|x_{i}\|$:
$\lambda_{\min}$ $(S_{ik})\|x_{k}\|^{2}$ (38)
$\leq\left[\sqrt{\max_{k\leq j<i}\mu_{j}}+\sqrt{\sum_{j=k}^{i-1}\mu_{j}B_{jk}}\right](\|x_{k}\|^{2}-\|x_{i}\|^{2}).$
Since $x_{k}$ is arbitrary, we may bound the norm of the state transition matrix as follows.
Theorem 5
Under the preceding notations, the following inequality holds:
$\|\Phi(N,k)\|^{2}\leq 1-\frac{\lambda_{\min}(S_{Nk})}{\left(\sqrt{\max_{k\leq j<N}\mu_{j}}+\sqrt{\sum_{j=k}^{N-1}\mu_{j}B_{jk}}\right)^{2}}.$ (39)
A key result in mathematical analysis is that, let $0\leq a_{k}<1$, then
$\prod_{i=1}^{\infty}(1-a_{i})=0\Leftrightarrow\sum_{i=1}^{\infty}a_{i}=\infty.$ (40)
Remark 8
Inspired by this fact, Guo set out to express the convergence of high-dimensional matrices in terms of divergence of one-dimensional series, which became the focus of his early work.
In our new framework, this basic fact yields an important corollary.
Corollary 1
Under the preceding notations, we have $\Phi(n,0)\to 0$ if
$\sum_{k=1}^{\infty}\frac{\lambda_{\min}(S_{t_{k}t_{k-1}})}{\left(\sqrt{\max_{t_{k-1}\leq j<t_{k}}\mu_{j}}+\sqrt{\sum_{j=t_{k-1}}^{t_{k}-1}\mu_{j}B_{jt_{k-1}}}\right)^{2}}=\infty.$ (41)
where $t_{k}$ is a strictly increasing natural number sequence which tends to infinity, i.e. $t_{k}\to\infty$. In particular, the SG algorithm is strongly consistent.
Remark 9
This result relies on a “time-scale rescaling” technique, which is not only natural but essential. The core idea is to partition the time axis into intervals $t_{k-1},t_{k})$ and analyze the system over these aggregated blocks. This approach is necessary because if we considered each matrix individually (i.e., interval length 1), the minimum eigenvalue $\lambda_{\min}(S_{t_{k}t_{k-1}})$ would be zero due to rank deficiency, causing the criterion to fail.
The weight sequence $\mu_{j}$ can be chosen as a streaming statistic of the regressors $\phi_{k}$’s, meaning each $\mu_{j}$ is computed from $\phi_{j}$ and $\mu_{j-1}$ alone. This property is particularly useful for online computation.
4 Convergence Analysis
In this section, we substitute $\phi_{j}$ with $\varphi_{j}/\sqrt{r_{j}}$, yielding the following estimate by Cauchy-Schwarz inequality:
$B_{jk}$ $=\sum_{l=k}^{j-1}(\phi_{j}^{\top}\phi_{l})^{2}=\sum_{l=k}^{j-1}\frac{(\varphi_{j}^{\top}\varphi_{l})^{2}}{r_{j}r_{l}}$ (42)
$\leq\sum_{l=k}^{j-1}\frac{\|\varphi_{j}\|^{2}}{r_{j}}\frac{\|\varphi_{l}\|^{2}}{r_{l}}=\frac{\|\varphi_{j}\|^{2}}{r_{j}}\sum_{l=k}^{j-1}\frac{r_{l}-r_{l-1}}{r_{l}}$
$\leq\frac{\|\varphi_{j}\|^{2}}{r_{j}}\int_{r_{k-1}}^{r_{j-1}}\frac{dx}{x}=\frac{r_{j}-r_{j-1}}{r_{j}}\log\frac{r_{j-1}}{r_{k-1}}.$
Substituting this into the denominator of (41), and choosing $\mu_{j}=r_{j}$, we have
$\sum_{j=t_{k-1}}^{t_{k}-1}\mu_{j}B_{jt_{k-1}}\leq\sum_{j=t_{k-1}}^{t_{k}-1}(r_{j}-r_{j-1})\log\frac{r_{j-1}}{r_{t_{k-1}-1}}$ (43)
$\leq\int_{r_{t_{k-1}-1}}^{r_{t_{k}-1}}\log xdx-(r_{t_{k}-1}-r_{t_{k-1}-1})\log r_{t_{k-1}-1}$
$=r_{t_{k}-1}\log r_{t_{k}-1}-r_{t_{k}-1}-r_{t_{k-1}-1}\log r_{t_{k-1}-1}$
$\quad+r_{t_{k-1}-1}-(r_{t_{k}-1}-r_{t_{k-1}-1})\log r_{t_{k-1}-1}$
$=r_{t_{k}-1}\left(\log r_{t_{k}-1}-\log r_{t_{k-1}-1}\right)-(r_{t_{k}-1}-r_{t_{k-1}-1}).$
This yields a applicable sufficient condition for the strong consistency.
Corollary 2
Under the preceding notations, we have $\Phi(n,0)\to 0$ if
$\sum_{k=1}^{\infty}\frac{\lambda_{\min}(S_{t_{k}t_{k-1}})}{D_{k}}=\infty.$ (44)
where $t_{k}$ is defined similarly, $S_{t_{k}t_{k-1}}=\sum_{i=t_{k-1}}^{t_{k}-1}\varphi_{i}\varphi_{i}^{\top}$, and $D_{k}=r_{t_{k}-1}\left(\log r_{t_{k}-1}-\log r_{t_{k-1}-1}\right)-(r_{t_{k}-1}-r_{t_{k-1}-1})$.
This lemma serves a function analogous to the ‘time inverse function m(t)m(t)” introduced by Chen and Guo, but admits a formulation that facilitates the subsequent analysis.
Lemma 1
Assume Assumption 3 is satisfied, then there exists strictly increasing natural number sequence $\{t_{k}\}$, such that
$\frac{k}{l}<\frac{r_{t_{k}}}{r_{t_{k-1}}}<lk,$ (45)
i.e., $r_{t_{k}}=\Theta(kr_{t_{k-1}})$.
{pf}
Define $t_{k}=\min\{j:r_{j}\geq k!\}$. By definition and Assumption 3,
$k!\leq r_{t_{k}}\leq lr_{t_{k}-1}<lk!$ (46)
Substituting $k$ by k1k-1 gives
$(k-1)!\leq r_{t_{k-1}}<l(k-1)!$ (47)
Combining these yields the desired inequality.
5 Main Theorem
Now we are fully prepared to prove the main theorem.
Theorem 6
Assume Assumption 3 is satisfied, and condition number growth
$\kappa\left(\sum_{i=1}^{n}\varphi_{i}\varphi_{i}^{\top}\right)=O((\log r_{n})^{\alpha}),$ (48)
where $\alpha<1$, then we have $\Phi(n,0)\to 0$.
{pf}
For convenience, set
$S_{n}=\sum_{i=1}^{n}\varphi_{i}\varphi_{i}^{\top}.$ (49)
The matrix in (44) , $S_{t_{k}t_{k-1}}$, can be expressed as the difference of two cumulative matrices:
$S_{t_{k}t_{k-1}}=S_{t_{k}-1}-S_{t_{k-1}-1}.$ (50)
We use Weyl’s inequality for the eigenvalues of a sum of Hermitian matrices, which states that
$\lambda_{\min}(S_{t_{k}t_{k-1}})\geq\lambda_{\min}(S_{t_{k}-1})-\lambda_{\max}(S_{t_{k-1}-1}).$ (51)
Now, we find a lower bound for the first term $\lambda_{\min}(S_{t_{k}-1})$ and an upper bound for the second term $\lambda_{\max}(S_{t_{k-1}-1})$. A lower bound is derived as follows:
$\lambda_{\min}(S_{t_{k}-1})$ $\geq\frac{\lambda_{\max}(S_{t_{k}-1})}{M(\log r_{t_{k}-1})^{\alpha}}$ (52)
$\geq\frac{1}{d}\frac{\mathrm{tr}(S_{t_{k}-1})}{M(\log r_{t_{k}-1})^{\alpha}}=\frac{1}{d}\frac{r_{t_{k}-1}-1}{M(\log r_{t_{k}-1})^{\alpha}}.$
The upper bound comes from monotonicity of $r_{n}$:
$\lambda_{\max}(S_{t_{k-1}-1})\leq\mathrm{tr}(S_{t_{k-1}-1})=r_{t_{k-1}-1}-1.$ (53)
Substituting the results back into the Weyl’s inequality gives the final lower bound:
$\lambda_{\min}(S_{t_{k}t_{k-1}})\geq\frac{1}{d}\frac{r_{t_{k}-1}-1}{M(\log r_{t_{k}-1})^{\alpha}}-(r_{t_{k-1}-1}-1).$ (54)
Choose $t_{k}$ as in Lemma 1. By Stirling’s formula, we have
$\log r_{t_{k}-1}\leq\log(lk!)\lesssim k\log k.$ (55)
This offers a lower bound of $\lambda_{\min}(S_{t_{k}t_{k-1}})$ in terms of $k$:
$\lambda_{\min}(S_{t_{k}t_{k-1}})$ $\gtrsim\frac{k!}{(k\log k)^{\alpha}}-(k-1)!$ (56)
$=(k-1)!\left(\frac{k^{1-\alpha}}{(\log k)^{\alpha}}-1\right)\gtrsim\frac{k!}{(k\log k)^{\alpha}}.$
Besides, note that
$D_{k}$ $\leq r_{t_{k}-1}\left(\log r_{t_{k}-1}-\log r_{t_{k-1}-1}\right)$ (57)
$\lesssim r_{t_{k}}\log\frac{r_{t_{k}}}{r_{t_{k-1}}}<r_{t_{k}}\log lk\lesssim k!\log k.$
To estimate the summand in (44), we obtain
$\frac{\lambda_{\min}(S_{t_{k}t_{k-1}})}{D_{k}}\gtrsim\frac{k!}{(k\log k)^{\alpha}}\frac{1}{k!\log k}=\frac{1}{k^{\alpha}(\log k)^{1+\alpha}}.$ (58)
Since the series on the right-hand side diverges, it follows that $\Phi(n,0)\to 0$.
6 Conclusions
This study has resolved a long-standing conjecture by Chen and Guo, demonstrating that the stochastic gradient algorithm achieves strong consistency even under non-persistent excitation, provided the condition number of the Fisher information matrix grows as $O((\log r_{n})^{\alpha})$ for any $0\leq\alpha<1$. Our work has significantly broadened the known sufficient condition from $\alpha\leq 1/3$ and has established $\alpha=1$ as the critical threshold. By introducing a novel algebraic framework, we have provided sharper matrix bounds and a more transparent proof. This advancement not only deepens the theoretical understanding of stochastic gradient optimization methods but also challenges researchers to finally settle the boundary case of $\alpha=1$. Future work should extend this framework to other step-size rules and validate its implications in practical deep learning and adaptive control scenarios.
{ack}
The authors are deeply grateful to Professor Lei Guo (IFAC Fellow, IEEE Fellow, AMSS), the first author’s supervisor, for his invaluable guidance and for posing the fundamental research question that motivated this work. His insightful comments and continuous support have greatly contributed to this study. The authors are also indebted to Yujing Liu, Cheng Zhao and Xin Zheng for their valuable feedback and stimulating discussions.